I need to determine the slope and y-intercept of the following equation
2x+3y=5
I believe the slope is 2x and that the y intercept is 3y but I am not sure.
Thank you

Answer by rfer(16322)

put in slope intercept form first
y=mx+b
3y=-2x+5
y=-2/3x+5/3
-----------
slope m=-2/3
y int =5/3 (0,5/3)